OpenStudy (anonymous):

A yogurt shop sells yogurt in small, medium, and large cups. There are 8 different flavors and 4 different toppings from which to choose. How many different yogurt selections are possible?

OpenStudy (anonymous):

multiply all the choices together i .e. compute \[3\times 8\times 4\]

OpenStudy (anonymous):

that is called the "counting principle"

OpenStudy (anonymous):

if there are \(m\) ways to do one thing and \(n\) ways to do another, then the number of ways you can do them together is \(m\times n\)